# FedALT: Federated Fine-Tuning through Adaptive Local Training with Rest-of-the-World LoRA

### v1

@@ Abstract
Fine-tuning large language models (LLMs) in federated settings enables privacy-preserving adaptation but suffers from cross-client interference due to model aggregation. Existing federated LoRA fine-tuning methods, primarily based on FedAvg, struggle with data heterogeneity, leading to harmful cross-client interference and suboptimal personalization. In this work, we propose FedALT, a novel personalized federated LoRA fine-tuning algorithm that fundamentally departs from FedAvg. Instead of using an aggregated model to initialize local training, each client continues training its individual LoRA while incorporating shared knowledge through a separate Rest-of-the-World (RoTW) LoRA component. To effectively balance local adaptation and global information, FedALT introduces an adaptive mixer that dynamically learns input-specific weightings between the individual and RoTW LoRA components using the Mixture-of-Experts (MoE) principle. Through extensive experiments on NLP benchmarks, we demonstrate that FedALT significantly outperforms state-of-the-art personalized federated LoRA fine-tuning methods, achieving superior local adaptation without sacrificing computational efficiency.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Illustration of FedALT. Instead of directly aggregating local LoRA modules from each client using FedAvg, FedALT introduces a frozen Rest-of-World LoRA component to transmit shared global knowledge while preserving client-specific adaptations through Individual LoRA modules. The adaptive mixer dynamically combines the RoTW LoRA and Individual LoRA. See details in Section 3.
@@
Large language models (LLMs) (Kenton & Toutanova, 2019; Brown et al., 2020; Raffel et al., 2020; Touvron et al., 2023a; Zhou et al., 2024; Zeng et al., 2022; Touvron et al., 2023b) have demonstrated exceptional capabilities in language understanding and generation, enabling a wide range of applications. Adapting pretrained LLMs to specialized domains or further enhancing their performance requires fine-tuning with task-specific data (Minaee et al., 2024). However, full fine-tuning, which involves updating all model parameters, is computationally expensive and often impractical for real-world deployment. To address this limitation, several parameter-efficient fine-tuning (PEFT) (Ding et al., 2023; Fu et al., 2023; Han et al., 2024; Liu et al., 2022; Li & Liang, 2021; Lester et al., 2021) methods have been proposed, with low-rank adaptation (LoRA) (Hu et al., 2021) being one of the most widely used. LoRA reduces the number of trainable parameters by integrating low-rank matrices into the model, significantly reducing computational costs.
Fine-tuning LLMs requires substantial data to adapt to downstream tasks, yet this data is often distributed across multiple institutions where privacy regulations and security concerns prohibit direct sharing. Federated learning (FL) (McMahan et al., 2017) enables collaborative fine-tuning without exposing raw data, making it an essential framework for privacy-preserving adaptation. However, most existing federated LoRA fine-tuning approaches (Zhang et al., 2024; Sun et al., 2024; Wang et al., 2024d; Guo et al., 2024; Bian et al., 2024b) assume a single global model, overlooking the highly heterogeneous data distributions among clients. A one-size-fits-all model often results in suboptimal performance and fairness issues, especially in real-world scenarios where client data varies significantly in size, domain, and distribution. For example, in healthcare (Kim et al., 2024), hospitals may require fine-tuned models tailored to specific patient demographics or disease profiles. Without personalization, federated fine-tuning fails to capture such variations, leading to poor local adaptation and inefficient resource utilization. Despite the critical need for personalized solutions, personalized federated LoRA fine-tuning (Yang et al., 2024a; Anonymous, 2025; Qi et al., 2024) remains largely underexplored.
@@
Federated Averaging (FedAvg) (McMahan et al., 2017) is perhaps the most widely used FL framework, serving as the foundation for numerous follow-up works that introduce various optimizations. Its core principle is to aggregate locally trained client models into a global model, which then initializes the next round of local training. Most personalized FL algorithms (Yang et al., 2024a; Qi et al., 2024; Tan et al., 2022; Deng et al., 2020; Collins et al., 2021; Fallah et al., 2020), despite their implementation differences, build upon the FedAvg framework by incorporating personalization mechanisms—and personalized federated LoRA fine-tuning follows the same paradigm. For example, the recently proposed FedDPA (Yang et al., 2024a) algorithm personalizes LoRA by introducing both global and local LoRA components. The global LoRA component is trained using FedAvg, while the local LoRA component is trained independently on each client’s private data. This local adaptation can occur either after the global LoRA stabilizes or alternatedly with global LoRA training, in which case the final output is computed as a weighted combination of both components using a pre-defined weighting factor. However, while personalization improves the average performance across clients, some clients may experience degraded performance, even falling behind simple local fine-tuning—that is, not participating in FL at all. This issue, also observed in our motivational study in Section 2.3, raises a critical question: Is FedAvg truly the most effective framework for developing personalized federated LoRA fine-tuning methods in heterogeneous settings?
In this paper, we propose a novel personalized federated LoRA fine-tuning algorithm that departs fundamentally from the FedAvg paradigm. Our method is motivated by identifying two key limitations of FedAvg-based approaches, such as FedDPA, in handling FL challenges under heterogeneous client settings. First, FedAvg-based approaches suffer from harmful cross-client interference due to their reliance on a global model aggregated from client models. This aggregation step, intended to capture shared knowledge, often disrupts individual client adaptations when data distributions are substantially different. As a result, any improvements achieved during local fine-tuning can be negated by aggregation, making it ineffective for clients whose distributions deviate significantly from the global model. For such clients, initializing the next training round with the aggregated model may offer little to no benefit—or even degrade performance. Second, FedAvg-based approaches lack an effective mechanism to balance local and global information in personalization. For example, FedDPA combines global and local LoRA components using a pre-defined weighting factor during training. However, improper weighting can significantly harm personalized fine-tuning, particularly when the global model is assigned excessive weight. This issue is especially problematic for clients where the global model is poorly suited, leading to suboptimal performance despite the presence of a local adaptation mechanism.
Overall, the FedAvg paradigm structures the personalization mechanisms around the aggregated model, treating local training as a supplement to the global model. In contrast, our approach emphasizes the role of local training in heterogeneous settings, focusing on personalization by incorporating useful global information directly into local training. Specifically, we eliminate the key FedAvg step of using the aggregated model to initialize each local training round. Instead, clients continue training on their own local models from the previous round. Information from other clients’ models is aggregated into a “rest-of-the-world” (RoTW) LoRA component, which supports local training but remains frozen during each round. Furthermore, to effectively leverage the RoTW component, we introduce an “adaptive mixer” that dynamically learns the optimal weighting between the individual LoRA and RoTW LoRA components using the Mixture-of-Experts (MoE) principle (Jordan & Jacobs, 1994). This adaptive weighting enables the model to selectively integrate relevant global knowledge while maintaining local model adaptation. We name our approach FedALT, short for Federated Fine-tuning through Adaptive Local Training with Rest-of-the-World LoRA.
We highlight that FedALT is particularly well-suited for federated LoRA fine-tuning due to its low computational complexity and minimal memory requirements. Although a separate RoTW LoRA is maintained alongside each client’s individual LoRA component, the RoTW LoRA is not trained during each round. Furthermore, LoRA modules are significantly smaller than the full model, meaning the added memory overhead is negligible. Through comprehensive experiments on heterogeneous FL benchmarks across various NLP tasks, we demonstrate that FedALT outperforms state-of-the-art approaches by a significant margin.
@@ 2.2 Heterogeneous Federated Fine-Tuning
$\min_{\mathcal{W}}\frac{1}{K}\sum_{k=1}^{K}L(X_{k},Y_{k},\mathbf{W}_{k}),$ (2)
where $L(\cdot)$ represents the loss function, and $\mathcal{W}$ denotes the set of personalized fine-tuned models $\{\mathbf{W}_{k}\}_{k=1}^{K}$.
@@ 2.3 Motivational Study
Federated fine-tuning operates in highly heterogeneous environments where client datasets can differ significantly. To understand the effects of such heterogeneity, we conducted a motivational study that highlights both the benefits and challenges of applying FL in these settings.
Figure 2: Motivational study results shows that while FedIT underperforms Local Only on some clients due to harmful cross-client interference. FedIT-split, using multiple smaller LoRAs, fails to mitigate this interference, highlighting the limitations of the FedAvg paradigm. See details in Section 2.3.
@@
In FL settings, the data collected by each client can be highly diverse (Huang et al., 2022; Mendieta et al., 2022; Wang et al., 2024b). We conducted a motivational experiment using the Flan dataset (Chung et al., 2024), selecting 8 tasks and creating 8 clients, each specializing in a single task. To demonstrate potential interference in federated fine-tuning, we compared two baseline methods. The first method, FedIT (Zhang et al., 2024), combines FedAvg with LoRA fine-tuning. The second method is Local Only LoRA fine-tuning, where each client fine-tunes its LoRA locally without any communication or information sharing between clients. The results, shown in Figure 2, support our first observation. On one hand, federated fine-tuning using FedIT achieves better performance than local fine-tuning on tasks Coreference Resolution and Paraphrase Detection, confirming that FL can enhance learning by enabling clients to benefit from shared knowledge. On the other hand, for tasks Commonsense Reasoning and Text Classification, FedIT underperforms compared to individual local fine-tuning. This dual outcome underscores both the promise and the pitfalls of federated fine-tuning in heterogeneous settings.
@@
Having established the existence of harmful interference, we next explored potential mitigation strategies. In centralized LoRA fine-tuning, HydraLoRA (Tian et al., 2024) found that splitting a single large LoRA into multiple smaller LoRAs reduces cross-domain interference. However, FL introduces additional challenges, as data domains are distributed across clients. To test whether multiple smaller LoRAs could mitigate interference in FL, we implemented FedAvg with multiple parallel LoRAs (i.e. FedIT-split). However, as shown in Figure 2, its performance remained comparable to FedIT (FedAvg with a single large LoRA), indicating that interference persisted. This result is expected: in centralized settings, interference occurs within a shared dataset and can be mitigated by partitioning LoRAs. In FL, interference arises primarily during server aggregation, and simply introducing multiple LoRAs does not prevent the aggregated model from disrupting local adaptations. Consequently, direct adaptations of centralized interference-mitigation strategies fail to address the core issue in FL.
@@ 3 Proposed Method: FedALT
Our motivational studies reveal that harmful interference in federated learning arises from server aggregation and the use of the aggregated model to initialize subsequent training rounds. In the FedAvg paradigm, a global LoRA serves as a shared initialization point, allowing clients to fine-tune locally on their respective datasets. After fine-tuning, the updated LoRAs are averaged on the server and redistributed to clients as the new initialization. However, when client data distributions differ significantly, aggregation disrupts individual adaptations, often negating improvements from local training and hindering personalization.
Can cross-client interference be mitigated? The simplest and most direct solution is local fine-tuning using only a client’s own dataset. This approach eliminates interference entirely but fails to leverage shared knowledge from other clients. To address this trade-off, we propose a novel personalized federated LoRA fine-tuning algorithm that fundamentally diverges from the FedAvg paradigm, prioritizing local training while selectively integrating useful global information.
Our method introduces two key innovations: 1. Each client maintains two LoRA modules: an Individual LoRA, which captures locally learned information and is updated exclusively during the client’s fine-tuning, and a Rest-of-the-World (RoTW) LoRA, which aggregates information from all other clients. Crucially, the RoTW LoRA remains fixed during local fine-tuning, ensuring that aggregation does not interfere with a client’s adaptation. 2. An adaptive mixer dynamically learns the optimal input-specific weighting between the Individual LoRA and RoTW LoRA components, leveraging the Mixture-of-Experts (MoE) principle to balance personalization and global knowledge integration.
In the following of this section, we detail the design and functionality of these components.
3.1 Individual and Rest-of-the-World (RoTW) LoRA
Each client $k$ maintains two types of LoRAs: an Individual LoRA, denoted by $\mathbf{A}^{L}_{k}/\mathbf{B}^{L}_{k}$, and a Rest-of-the-World (RoTW) LoRA, denoted by $\mathbf{A}^{R}_{k}/\mathbf{B}^{R}_{k}$. The Individual LoRA captures information obtained locally by the client and is isolated from the information of other clients. In contrast, the RoTW LoRA represents the average of the Individual LoRAs from all other clients and remains frozen during the client’s local fine-tuning. Specifically, at the end of each round, client $k$ updates its Individual LoRA $\mathbf{A}^{L}_{k}/\mathbf{B}^{L}_{k}$ through local training. The RoTW LoRA for client $k$ is then computed as:
@@ 3.1 Individual and Rest-of-the-World (RoTW) LoRA
After clients upload their local LoRAs to the server, the server computes the RoTW LoRA for each client individually and distributes it back to them. Alternatively, the server can calculate the global average of all clients’ individual LoRAs and distribute this average back to the clients. Each client then determines its RoTW LoRA by subtracting its individual LoRA from the global average LoRA.
Now, consider the scenario where a weight $\frac{1}{K}$ is assigned to the local LoRA and $1-\frac{1}{K}$ to the RoTW LoRA when combining their outputs. In this case, the combined LoRA is mathematically equivalent to the averaged LoRA. This equivalence arises naturally from the definitions of the RoTW and averaged LoRAs. If these two LoRAs are trained simultaneously on local datasets, the approach essentially reduces to the standard FedAvg (or FedIT) algorithm, as all clients would end up with the same combined LoRA.
However, it is crucial to emphasize that the RoTW LoRA is not further updated on clients using their local datasets. Instead, it is only updated through averaging at the end of each learning round. Training the RoTW LoRA on local datasets is unnecessary and potentially counterproductive, as subsequent averaging could cancel out any improvements. Furthermore, skipping local training for the RoTW LoRA reduces the computational workload on clients by half.
@@ 3.2 Dynamic Mixture-of-Experts Mechanism
A natural question arises: why not simply add the RoTW LoRA directly to the pre-trained model, i.e., $\mathbf{W}_{0}\leftarrow\mathbf{W}_{0}+\mathbf{B}^{R}_{k}\mathbf{A}^{R}_{k}$, since the RoTW LoRA remains fixed during local training? This approach would allow each client to maintain only a single Individual LoRA, saving memory space. However, there are two key reasons why this is not desirable:
@@
    Potential Contamination of the Pre-trained Model: If the RoTW LoRA underperforms, adding it directly to the pre-trained model risks contaminating it, making further corrections difficult.
@@
    Loss of Flexibility: Directly incorporating the RoTW LoRA enforces a fixed weight, preventing dynamic adjustment between the Individual LoRA and shared information for different inputs. This lack of flexibility is problematic because different inputs may benefit variably from the local model versus the globally averaged model. A “one-size-fits-all” approach often fails to perform optimally across diverse inputs.
To address these challenges, we propose a dynamic weighting mechanism for combining the Individual LoRA and the RoTW LoRA, enabling input-specific flexibility. Specifically, we introduce a mixer to dynamically adjust the contributions of the two LoRAs for each input. During the forward pass of each local epoch, the model’s output is computed as:
@@
where $\alpha_{k}(x)$ determines for client $k$ the contribution weight of the Individual LoRA, and $1-\alpha_{k}(x)$ determines the contribution of the RoTW LoRA.
Given the complexity of modern LLMs, where LoRA modules are inserted at multiple layers, learning $\alpha(x)$ is non-trivial. Inspired by (Jordan & Jacobs, 1994), we adopt a Mixture-of-Experts (MoE) approach to compute $\alpha(x)$. Specifically, we introduce a mixer $\mathbf{G}_{k}\in\mathbb{R}^{2\times d}$ which is a dense layer with trainable weights (a transformation matrix), followed by a softmax function, to learn the weighting:

### v2

# FedALT: Federated Fine-Tuning through Adaptive Local Training with Rest-of-World LoRA

@@ Abstract
Fine-tuning large language models (LLMs) in federated settings enables privacy-preserving adaptation but suffers from cross-client interference due to model aggregation. Existing federated LoRA fine-tuning methods, primarily based on FedAvg, struggle with data heterogeneity, leading to harmful cross-client interference and suboptimal personalization. In this work, we propose FedALT, a novel personalized federated LoRA fine-tuning algorithm that fundamentally departs from FedAvg. Instead of using an aggregated model to initialize local training, each client continues training its individual LoRA while incorporating shared knowledge through a separate Rest-of-World (RoW) LoRA component. To effectively balance local adaptation and global information, FedALT introduces an adaptive mixer that dynamically learns input-specific weightings between the individual and RoW LoRA components, drawing conceptual foundations from the Mixture-of-Experts (MoE) paradigm. Through extensive experiments on NLP benchmarks, we demonstrate that FedALT significantly outperforms state-of-the-art personalized federated LoRA fine-tuning methods, achieving superior local adaptation without sacrificing computational efficiency.
@@ 1 Introduction
Figure 1: Illustration of FedALT. Instead of directly aggregating local LoRA modules from each client using FedAvg, FedALT introduces a frozen RoW LoRA component to transmit shared global knowledge while preserving client-specific adaptations through Individual LoRA. The adaptive mixer dynamically combines the RoW LoRA and Individual LoRA. See details in section 3.
@@
Large language models (LLMs) [19, 4, 31, 37, 50, 48, 38] have demonstrated exceptional capabilities in language understanding and generation, enabling a wide range of applications. Adapting pretrained LLMs to specialized domains or further enhancing their performance requires fine-tuning with task-specific data [29]. However, full fine-tuning, which involves updating all model parameters, is computationally expensive and often impractical for real-world deployment. To address this limitation, several parameter-efficient fine-tuning (PEFT) [9, 11, 14, 25, 24, 23] methods have been proposed, with low-rank adaptation (LoRA) [16] being one of the most widely used. LoRA reduces the number of trainable parameters by integrating low-rank matrices into the model, significantly reducing computational costs.
Fine-tuning LLMs necessitates substantial data volumes for adaptation to downstream tasks; however, this requisite data often resides across multiple institutions where privacy regulations and security protocols prohibit direct data sharing. Federated Learning (FL) [27] facilitates collaborative fine-tuning without raw data exposure, establishing itself as an essential paradigm for privacy-preserving adaptation. Nonetheless, existing federated LoRA fine-tuning methodologies [49, 34, 43, 12, 3] predominantly presuppose a singular global model, neglecting the inherent data heterogeneity among participating clients. Such a homogeneous modeling approach frequently yields suboptimal performance and introduces fairness disparities, particularly in real-world deployments where client data exhibits significant variance in volume, domain, and distributional characteristics. While limited research on personalized federated LoRA fine-tuning [15, 30] exists, these approaches primarily address label heterogeneity cases. However, in realistic LLM fine-tuning contexts, clients’ learning objectives naturally diverge due to diverse business imperatives [47]. Personalized federated LoRA fine-tuning addressing more realistic task-heterogeneous scenarios [45] remains insufficiently investigated in the current literature.
@@
Federated Averaging (FedAvg) [27] is perhaps the most widely used FL framework, serving as the foundation for numerous follow-up works that introduce various optimizations. Its core principle is to aggregate locally trained client models into a global model, which then initializes the next round of local training. Most personalized FL algorithms [45, 30, 35, 8, 7, 10], despite their implementation differences, build upon the FedAvg framework by incorporating personalization mechanisms—and personalized federated LoRA fine-tuning follows the same paradigm. For example, the recently proposed FedDPA [45] algorithm personalizes LoRA by introducing both global and local LoRA components. The global LoRA component is trained using FedAvg, while the local LoRA component is trained independently on each client’s private data. This local adaptation can occur either after the global LoRA stabilizes or alternatedly with global LoRA training, in which case the final output is computed as a weighted combination of both components using a pre-defined weighting factor. However, our motivational study (see section 2.3) reveals that FedAvg-based approaches may degrade client-specific performance due to harmful cross-client interference. We identify two fundamental limitations of FedAvg-based methods like FedDPA in heterogeneous FL settings. First, they suffer from harmful cross-client interference due to reliance on global aggregation, which can negate individual client improvements achieved during local fine-tuning. Second, these methods lack an effective mechanism for balancing global and local information, relying on fixed or predefined weighting factors, which can be suboptimal for clients poorly represented by the global model.
To address these critical limitations, we propose a novel personalized federated LoRA fine-tuning approach called FedALT (short for Federated Fine-tuning through Adaptive Local Training with Rest-of-World LoRA.). Distinctly different from FedAvg-based methods, FedALT does not aggregate local models to initialize training rounds. Instead, each client continues training on their own previously trained local models, supplemented by a frozen Rest-of-World (RoW) LoRA component. The RoW LoRA captures global knowledge without updating during local training rounds, thus avoiding interference with client-specific adaptations. To leverage RoW LoRA effectively, we introduce an adaptive mixer, inspired by Mixture-of-Experts (MoE) [18], which dynamically and optimally combines global and local information for each client’s data sample. Our contributions are as follows:
  • We highlight that personalized federated fine-tuning for clients with heterogeneous tasks has been insufficiently explored and identify critical limitations in existing methods that build upon the FedAvg paradigm.
  • We propose FedALT, a novel personalization framework featuring a frozen RoW LoRA and an adaptive mixer, effectively addressing these limitations by dynamically balancing global and local model adaptations.
  • We perform comprehensive experiments on two widely-used LLMs (Bloom [21] and Llama 2 [38]) using the Flan benchmark [6], demonstrating that FedALT significantly outperforms existing federated LoRA fine-tuning methods.
@@ 2.2 Heterogeneous Federated Fine-Tuning
$\min_{\mathcal{W}}\frac{1}{K}\sum_{k=1}^{K}L_{k}(X_{k},Y_{k},\mathbf{W}_{k}),$ (2)
where $L_{k}(\cdot)$ represents the loss function for client $k$, and $\mathcal{W}$ denotes the set of personalized fine-tuned models $\{\mathbf{W}_{k}\}_{k=1}^{K}$.
@@ 2.3 Motivational Study
Federated fine-tuning operates in highly heterogeneous environments where client tasks can differ significantly. To understand the effects of such heterogeneity, we conducted a motivational study that highlights both the benefits and challenges of applying FL in these settings.
@@
In FL settings, the data collected by each client can be highly diverse [17, 28, 40]. We conducted a motivational experiment using the Flan dataset [6], selecting 8 tasks and creating 8 clients, each specializing in a single task. To demonstrate potential interference in federated fine-tuning, we compared two baseline methods. The first method, FedIT [49], combines FedAvg with LoRA fine-tuning. The second method is Local Only LoRA fine-tuning, where each client fine-tunes its LoRA locally without any communication or information sharing between clients.
Figure 2: Motivational study results. See details in section 2.3.
The results, shown in fig. 2, support our first observation. On one hand, federated fine-tuning using FedIT achieves better performance than local fine-tuning on tasks Coreference Resolution and Paraphrase Detection, confirming that FL can enhance learning by enabling clients to benefit from shared knowledge. On the other hand, for tasks Commonsense Reasoning and Text Classification, FedIT underperforms compared to individual local fine-tuning. This dual outcome underscores both the promise and the pitfalls of federated fine-tuning in heterogeneous settings.
@@
Having established the existence of harmful interference, we next explored potential mitigation strategies. In centralized LoRA fine-tuning, HydraLoRA [36] found that splitting a single large LoRA into multiple smaller LoRAs reduces cross-domain interference. However, FL introduces additional challenges, as data domains are distributed across clients. To test whether multiple smaller LoRAs could mitigate interference in FL, we implemented FedAvg with multiple parallel LoRAs (i.e. FedIT-split). However, as shown in fig. 2, its performance remained comparable to FedIT (FedAvg with a single large LoRA), indicating that interference persisted. This result is expected: in centralized settings, interference occurs within a shared dataset and can be mitigated by partitioning LoRAs. In FL, interference arises primarily during server aggregation, and simply introducing multiple LoRAs does not prevent the aggregated model from disrupting local adaptations. Consequently, direct adaptations of centralized interference-mitigation strategies fail to address the core issue in FL.
@@ 3 Proposed Method: FedALT
Our motivational studies reveal that harmful interference in FL arises from server aggregation and the use of the aggregated model to initialize subsequent training rounds. In the FedAvg paradigm, a global LoRA serves as a shared initialization point, allowing clients to fine-tune locally on their respective datasets. After fine-tuning, the updated LoRAs are averaged on the server and redistributed to clients as the new initialization. However, when client data differ significantly, aggregation disrupts individual adaptations, often negating improvements from local training and hindering personalization. Can cross-client interference be mitigated? The simplest and most direct solution is local fine-tuning using only a client’s own dataset. This approach eliminates interference entirely but fails to leverage shared knowledge from other clients. To address this trade-off, we propose a novel personalized federated LoRA fine-tuning algorithm that fundamentally diverges from the FedAvg paradigm, prioritizing local training while selectively integrating useful global information.
Our method introduces two key innovations: 1. Each client maintains two LoRA modules: an Individual LoRA, which captures locally learned information and is updated exclusively during the client’s fine-tuning, and a Rest-of-World (RoW) LoRA, which aggregates information from all other clients. Crucially, the RoW LoRA remains fixed during local fine-tuning, ensuring that aggregation does not interfere with a client’s adaptation. 2. An adaptive mixer dynamically learns the optimal input-specific weighting between the Individual LoRA and RoW LoRA components to balance personalization and global knowledge integration. In the following of this section, we detail the design and functionality of these components. Additionally, a theoretical convergence analysis of the proposed method is provided in the appendix F.
3.1 Individual and Rest-of-World (RoW) LoRA
Each client $k$ maintains two types of LoRAs: an Individual LoRA, denoted by $\mathbf{A}^{L}_{k}/\mathbf{B}^{L}_{k}$, and a Rest-of-World (RoW) LoRA, denoted by $\mathbf{A}^{R}_{k}/\mathbf{B}^{R}_{k}$. The Individual LoRA captures information obtained locally by the client and is isolated from the information of other clients. In contrast, the RoW LoRA represents the average of the Individual LoRAs from all other clients and remains frozen during the client’s local fine-tuning. Specifically, at the end of each round, client $k$ updates its Individual LoRA $\mathbf{A}^{L}_{k}/\mathbf{B}^{L}_{k}$ through local training. The RoW LoRA for client $k$ is then computed as:
@@ 3.1 Individual and Rest-of-World (RoW) LoRA
After clients upload their local LoRAs to the server, the server computes the RoW LoRA for each client individually and distributes it back to them. Alternatively, the server can calculate the global average of all clients’ individual LoRAs and distribute this average back to the clients. Each client then determines its RoW LoRA by subtracting its individual LoRA from the global average LoRA.
Now, consider the scenario where a weight $\frac{1}{K}$ is assigned to the local LoRA and $1-\frac{1}{K}$ to the RoW LoRA when combining their outputs. In this case, the combined LoRA is mathematically equivalent to the averaged LoRA. This equivalence arises naturally from the definitions of the RoW and averaged LoRAs. If these two LoRAs are trained simultaneously on local datasets, the approach essentially reduces to the standard FedAvg (or FedIT) algorithm, as all clients would end up with the same combined LoRA.
However, it is crucial to emphasize that the RoW LoRA is not further updated on clients using their local datasets. Instead, it is only updated through averaging at the end of each learning round. Training the RoW LoRA on local datasets is unnecessary and potentially counterproductive, as subsequent averaging could cancel out any improvements. Furthermore, skipping local training for the RoW LoRA reduces the computational workload on clients by half.
@@ 3.2 Dynamic Mixture-of-Experts Mechanism
A natural question arises: why not simply add the RoW LoRA directly to the pre-trained model, i.e., $\mathbf{W}_{0}\leftarrow\mathbf{W}_{0}+\mathbf{B}^{R}_{k}\mathbf{A}^{R}_{k}$, since the RoW LoRA remains fixed during local training? This approach would allow each client to maintain only a single Individual LoRA, saving memory space. However, there are two key reasons why this is not desirable:
@@
    Potential Contamination of the Pre-trained Model: If the RoW LoRA underperforms, adding it directly to the pre-trained model risks contaminating it, making further corrections difficult.
@@
    Loss of Flexibility: Directly incorporating the RoW LoRA enforces a fixed weight, preventing dynamic adjustment between the Individual LoRA and shared information for different inputs. This lack of flexibility is problematic because different inputs may benefit variably from the local model versus the globally averaged model. A “one-size-fits-all” approach often fails to perform optimally across diverse inputs.
To address these challenges, we propose a dynamic weighting mechanism for combining the Individual LoRA and the RoW LoRA, enabling input-specific flexibility. Specifically, we introduce a mixer to dynamically adjust the contributions of the two LoRAs for each input. During the forward pass of each local epoch, the model’s output is computed as:
@@
where $\alpha_{k}(x)$ determines for client $k$ the contribution weight of the Individual LoRA, and $1-\alpha_{k}(x)$ determines the contribution of the RoW LoRA.
Given the complexity of modern LLMs, where LoRA modules are inserted at multiple layers, learning $\alpha(x)$ is non-trivial. Inspired by [18], we adopt a Mixture-of-Experts (MoE) approach to compute $\alpha(x)$. Specifically, we introduce a mixer $\mathbf{G}_{k}\in\mathbb{R}^{2\times d}$ which is a dense layer with trainable weights (a transformation matrix), followed by a softmax function, to learn the weighting: